If a man climbs up a wall that is 2 m high and gains 1,220 J of potential energy, then the mass of the man is 62.24 kg.


true or false


1
Expert's answer
2021-02-04T02:33:46-0500
"PE=mgh,""m=\dfrac{PE}{gh}=\dfrac{1220\ J}{9.8\ \dfrac{m}{s^2}\cdot2.0\ m}=62.24\ kg."

Answer:

True.